OpenStudy (anonymous):

7th grade math help plzzz

OpenStudy (anonymous):

A recipe that makes 4 servings calls for 3/4 cup of apple juicejuice and 2/3 cup of cranberry juice. How many cups of juice are there per serving?

OpenStudy (ericanoel912):

17/12 x 1/4= 17/48 I got 1 7/24 cups, but that looks totally wrong.. sorry :(

OpenStudy (eliasperez191):

yeah he is right i types it on my phone and i got something else sorry thats wrong i also got 17/24 cups

OpenStudy (anonymous):

that's not one of my answer choices

OpenStudy (ericanoel912):

what are the choices and @eliasperez191 it's ERICA lol :)

OpenStudy (anonymous):

Alright so together we have 35+34=2720 per 3 servings...so we can just divide that by 3 to get the per 1 serving... 2720÷3 we flip the second number and change this to multiplication 2720×13=2760=920

OpenStudy (anonymous):

5/7 17/48 1 5/12 5/3

OpenStudy (texaschic101):

3/4 + 2/3 = total juice 9/12 + 8/12 = total juice 17/12 = total juice now set up a proportion.. 4/(17/12) = 1/x ---(4 serv to 17/12 cups of j = 1 serv to x cups of j cross multiply (4)(x) = (17/12)(1) 4x = 17/12 x = (17/12) / 4 x = 17/12 * 1/4 x = 17/48 cups Is that an answer choice ?

OpenStudy (anonymous):

no I just gave you the answer choices

OpenStudy (texaschic101):

yes it is.....17/48 cups

OpenStudy (eliasperez191):

ya its 17/48

OpenStudy (anonymous):

thanx

OpenStudy (texaschic101):

anytime :)

OpenStudy (eliasperez191):

yea if u need anything just ask